# Mass-loaded magnetic explosions in the context of Magnetar Giant Flares and Fast Radio Bursts

### v1

@@ Abstract
Context. Magnetar flares are highly energetic and rare events where intense X and $\gamma$-ray emission is released from strongly magnetised neutron stars and are accompanied by mass ejection from the neutron star. Fast radio bursts are short and intense pulses of coherent radio emission. Their large dispersion measures support an extragalactic origin. While their exact origin still remains elusive, a substantial number of models associates them with strong magnetic field and high-energy relativistic plasma found in the vicinity of magnetars. There is growing evidence that some fast radio bursts are associated to flare-type events from magnetars.
Aims. We aim to provide a set of configurations describing a relativistic, spherical, mass-loaded, magnetic explosion. We proceed by solving the equations of relativistic magnetohydrodynamics, for a system that expands while maintaining its internal equilibrium.
Methods. We employ a semi-analytical approach for the solution of the equations of relativistic magnetohydrodynamics. We assume self-similarity in time and radius, axial symmetry, and separation of variables in the spherical and polar angle coordinate. This allows the reduction of the problem to the solution of a set of ordinary differential equations.
Results. We find the interdependent relation between pressure, mass density, Lorentz factor, and magnetic field that determines the detailed properties of the solutions. There exists a dichotomy of solutions that correspond to higher and lower density and thermal pressure compared to the external one. For stronger toroidal magnetic fields the maximum expansion velocity, permitted, becomes smaller compared to weaker toroidal fields. For given ratio of toroidal to poloidal field, the inclusion of pressure and mass density leads to either higher expansion velocity if the density and pressure are lower in the regions where the magnetic flux is higher; whereas, if the pressure and mass density are higher in the regions where the magnetic flux is higher, the expansion velocity is smaller.
Conclusions. The classes of these solutions can be applied to magnetar giant flares and fast radio bursts. The ones corresponding to overdensities and higher pressure can be associated to magnetar flares, whereas the ones corresponding to underdensities can be relevant to fast radio bursts corresponding to magnetically dominated events with low mass loading.
@@ 1 Introduction
Magnetars are hosts of highly energetic events, releasing radiation in X-ray and $\gamma$-rays. The most energetic of these events are giant flares, that release an isotropic luminosity up to $10^{47}$ erg/s (Hurley et al., 1999; Palmer et al., 2005; Hurley et al., 2005). Such events have been observed from extra-galactic sources (Beniamini et al., 2025; Trigg et al., 2025) and have been associated to strong magnetised explosions (Thompson & Duncan, 1995) originating from magnetars where a major reconfiguration of the global magnetic field has occurred (Lyutikov, 2006). Magnetar flares also are loaded with mass entrained from the crust and the interior of the neutron star (Frail et al., 1999; Demidov & Lyubarsky, 2023; Patel et al., 2025). Thus, a complete view of giant flares, requires the modelling of a mass-loaded magnetised explosion, reaching relativistic velocities.
@@
The collection of an ample sample of fast radio bursts (FRBs), which are millisecond-duration transients of extragalactic origin, allows to explore their diverse phenomenology, which point to coherent emission from compact astrophysical sources (Lorimer et al., 2007; Petroff et al., 2019). Since their discovery, FRBs have challenged theoretical models, demanding mechanisms capable of releasing up to $10^{40}$ erg in radio waves over millisecond timescales, often with repetition, polarization, and spectral properties indicative of strong magnetized environments (Marcote et al., 2020). The growing evidence linking at least a subset of FRBs to magnetars — particularly the association of FRB 200428 with the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) — has reinforced the idea that magnetic energy release in highly magnetized neutron stars can power these events.
@@
Despite this association, the detailed physics linking magnetic activity to FRB emission remains uncertain. Proposed mechanisms span a broad range, from magnetospheric reconnection and crustal cracking to relativistic shocks in magnetar winds (e.g. Lyubarsky, 2014; Beloborodov, 2017; Metzger et al., 2019). Most models assume magnetically dominated outflows capable of generating coherent radio emission through synchrotron maser, shocks or antenna processes, and possibly a combination of those (Lyubarsky, 2021). Theoretical and observational developments suggest that magnetar magnetospheres may often be mass-loaded by plasma ejected during starquakes, bursts, or fallback episodes (Thompson & Duncan, 1995; Wadiasingh & Timokhin, 2019; Yuan et al., 2020). Moreover, the connection between magnetar bursts and flares and FRBs that has been confirmed recently (Bochenek et al., 2020; Tavani et al., 2021; Beniamini et al., 2025), stresses the necessity to consider mass loaded magnetic detonations that may be related to an event originating from the neutron star interior and propagating to the magnetosphere (Bransgrove et al., 2025). The presence of dense plasma in a strongly magnetised environment may substantially modify the dynamics of magnetic explosions, their ability to accelerate particles, and the conditions for coherent emission.
@@
In this work, we explore mass-loaded magnetic explosions as potential origins of magnetar giant flares and FRBs. Using simplified relativistic magnetohydrodynamic (MHD) models, we examine how the interplay between magnetic field, plasma mass loading, and self-similar expansion shapes the energetics and temporal properties of such events. The proposed structures can provide an appropriate environment that can lead to the generation of short, intense events with energetics and timescales consistent with observed FRBs.
@@
To model the coupled evolution of the magnetic field and plasma during these explosive events, we adopt the formalism developed in Gourgouliatos & Vlahakis (2010), which provides a self-consistent description of self-similar, co-expanding magnetized outflows. In this framework, the plasma and magnetic field evolve together under a homologous (Hubble-type) expansion, preserving the topology of the magnetic field while allowing for a time-dependent redistribution of electro-magnetic and kinetic energy.
@@
The underlying concept traces back to the Prendergast (2005) family of self-similar, uniformly-expanding, axisymmetric magnetic equilibria. These equilibria, later extended by Gourgouliatos & Lynden-Bell (2008) to include non-linear poloidal currents and toroidal fields, demonstrated that magnetic structures can expand coherently, a key insight for modelling transient magnetic explosions in compact objects. Building on this foundation, Gourgouliatos & Vlahakis (2010) formulated a relativistic configuration that couples the magnetic and velocity fields in a co-expanding MHD flow, enabling analytical and numerical exploration of outflows that evolve self-similarly while conserving magnetic helicity and flux. In these models the field originates from an appropriately twisted dipolar located at the origin. Due to the singularity at the origin the field is studied beyond an inner radius, that serves as a lower boundary condition.
@@
More recently, Barkov et al. (2022) applied similar ideas for numerical solutions of magnetic explosions in magnetically dominated and mass-loaded environments, demonstrating how magnetic energy release can drive relativistic outflows and shocks capable of producing bright, short-duration transients. In the same work, analytical solutions of force-free magnetic fields, in the absence of thermal pressure and inertia forces, were presented. These solutions do not diverge at the origin. It was found that the maximum expansion rate that can be attained by the expanding structure depends on the amount of twist of the magnetic field, with highly twisted magnetic fields not being able to expand at highly relativistic rates. In the absence of any thermal pressure and inertia, the pressure on the surface of the boundary of the spherical explosion is highly anisotropic due to the magnetic field structure, i.e. in the dipole case it is maximum on the equator and becomes zero on the axis of symmetry. Such a structure if embedded within a uniform pressure environment it will expand side-ways on the equator and will form a highly anisotropic explosion, while it will lose its internal equilibrium and coherence. The inclusion of mass density and thermal pressure leads to configurations in the form of large-scale plasmoids whose surface pressure, including the electromagnetic term is more isotropic and under certain conditions it can even become completely isotropic, with the surface magnetic field completely vanishing, as has been shown in the non-relativistic approach (Gourgouliatos et al., 2010). Such structures continue their uniform expansion if they are situated in a uniform density medium and maintain their internal equilibrium (Gourgouliatos & Lyutikov, 2012).
@@
In the present work, we extend these ideas by embedding mass-loaded magnetic explosions within the relativistically exapanding self-similar MHD formalism of Gourgouliatos & Vlahakis (2010). This allows us to follow consistently the co-evolution of magnetic fields and plasma during an explosive expansion, explore the interplay between magnetic field strength, mass and pressure and its effect on the expansion rate. This, will allow the identification of the physical conditions under which such events can generate coherent radio emission with energetics and timescales comparable to magnetar flares and FRBs.
@@
This paper is organized as follows. In section 2, we outline the physical and mathematical framework for mass-loaded magnetic explosions and describe the key ingredients of our model. Section 3 presents the solutions and the exploration of the parameter space. The results are discussed in section 4, focusing on energy release, plasma dynamics, and physical properties of the solution. In section 5, we discuss the implications for astrophysical sources with main focus on magnetar bursts and FRBs. We conclude in section 6 and outline the limitations of the current model and the prospects for future work.
@@ 2 Mathematical Formulation
Consider a fluid of rest mass density $\rho$, pressure $p$ containing a magnetic field $\bm{B}$. The electric and magnetic field obey Maxwell’s equations:
@@
where $j^{0}$ is the electric charge density, $\bm{j}$ the electric current density and $c$ the speed of light. We will further consider ideal-MHD, thus there is no electric field in the co-moving frame of the fluid. For an observer who sees the fluid moving at velocity $\bm{v}$, normalised to the speed of light, Ohm’s law becomes:
@@
The continuity equation reads:
@@
where $\Gamma$ is the Lorentz factor:
$\Gamma=\frac{1}{\sqrt{1-v^{2}}}$ (7)
The momentum equation is:
@@
where $p$ is the pressure and the relativistic specific enthalpy is given by the following expression:
@@
where $\gamma$ is the ratio of the specific heats. Furthermore, the entropy equation reads:
@@
In what follows we assume a relativistic degenerate gas where $\gamma=4/3$. We will further consider spherical coordinates $(r,\theta,\phi)$ and prescribe a velocity given by the following expression:
@@
This velocity corresponds to uniform expansion from the centre. Thus, each fluid element moves at constant velocity without any acceleration and the system remains implicitly at force balance.
It is possible to find self-similar, separable solutions using as main variables the polar angle $\theta$ and the expansion velocity $v$. The full details of this procedure are shown Prendergast (2005); Gourgouliatos & Lynden-Bell (2008); Gourgouliatos & Vlahakis (2010), here we summarise the basic steps for clarity and consistency.
We express the magnetic field in terms of two functions: one related to the poloidal field $\Psi(\theta,v)$ and another to the toroidal $T(\theta,v)$:
@@
Combining Ohm’s law, equation (5), with the momentum equation (8), and taking the toroidal component, one deduces that the poloidal and toroidal functions are related to each other through the following expression:
@@
The pressure and the density are related through the entropy equation (10), which is satisfied for:
@@
where $Q(v,\theta)$ is an arbitrary function of $v$ and $\theta$. Upon integration of the continuity equation (6), we obtain that:
@@
and combining this with equation (14), we find that the pressure has the following form:
@@
where the quantities $Q$ $\tilde{p}$, $\tilde{\rho}$, all depending on the separation variables $v$ and $\theta$, related to each other through the following expression:
@@
Next, we evaluate the current and charge densities, through Maxwell’s equations (2) and (4) and substitute them into the momentum equation (8). We find that $\tilde{p}=\tilde{p}(\Psi)$ and that the system should satisfy the following equation:
@@
$+\frac{v^{2}}{(1-v^{2})^{2}}H\frac{dH}{d\Psi}+16\pi^{3}\sin^{2}\theta\,\frac{v^{4}}{(1-v^{2})^{3}}\frac{d\tilde{p}}{d\Psi}=0$ (18)
@@
Upon integration, and considering the fact the $B_{\phi}$ component of the field must remain finite:
@@
with $c_{0}$ a non-negative constant, which governs the intensity of the toroidal field or equivalently the poloidal current. The pressure term becomes:
@@
In the above expression $\tilde{p}_{0}$ is the pressure on the axis and the spherical surface of the plasmoid; and it has to be a non-zero constant. There is no constraint on the sign of $c_{1}$, however, the combination of $\tilde{p}_{0}$ and $c_{1}$ must ensure that $\tilde{p}$ remains non-negative throughout the solution.
Finally, substituting the expressions from equation (19) into equation (18) we obtain an ordinary differential equation for $g(v)$:
@@
The solution of the above equation will provide the radial structure of the configuration.

### v2

# Magnetic plasmoid explosions in the context of magnetar giant flares and fast radio bursts

@@ Abstract
Context. Magnetar flares are highly energetic and rare events in which intense X and $\gamma$-ray emission is released from strongly magnetised neutron stars. The events are also accompanied by mass ejection from the neutron star. Fast radio bursts are short and intense pulses of coherent radio emission. Their large dispersion measures support an extragalactic origin. While their exact origin still remains elusive, a substantial number of models associates them with strong magnetic field and high-energy relativistic plasma found in the vicinity of magnetars. There is growing evidence that some fast radio bursts are associated with flare-type events from magnetars.
Aims. We provide a set of configurations describing a relativistic, spherical, magnetic plasmoid explosion. We proceed by solving the equations of relativistic magnetohydrodynamics for a system that expands while maintaining its internal equilibrium.
Methods. We employed a semi-analytical approach to solve the equations of relativistic magnetohydrodynamics. We assumed self-similarity in time and radius, axial symmetry, and separation of variables in the spherical and polar angle coordinate. This allowed us to reduce the problem to solving a set of ordinary differential equations.
Results. We find the interdependent relation between pressure, mass density, Lorentz factor, and magnetic field that determines the detailed properties of the solutions. A dichotomy of solutions exists that correspond to higher and lower density and thermal pressure compared to the external one. For stronger toroidal magnetic fields, the maximum permitted expansion velocity becomes lower than the weaker toroidal fields. For a given ratio of the toroidal to the poloidal field, the inclusion of pressure and mass density leads to either a higher expansion velocity when the density and pressure are lower in regions with a higher magnetic flux or to a lower expansion velocity when the pressure and mass density are higher in regions with a higher magnetic flux.
Conclusions. These solution classes can be applied to magnetar giant flares and fast radio bursts. Those that corresponding to overdensities and higher pressure can be associated with magnetar flares, and those corresponding to underdensities can be relevant to fast radio bursts that correspond to magnetically dominated events with low mass loading.
@@ 1 Introduction
Magnetars are hosts of highly energetic events that release radiation in X-ray and $\gamma$-rays. The most energetic of these events are giant flares, which release an isotropic luminosity up to $10^{47}$ erg/s (Hurley et al., 1999; Palmer et al., 2005; Hurley et al., 2005). These events have been observed from extra-galactic sources (Beniamini et al., 2025; Trigg et al., 2025) and have been associated with strong magnetised explosions (Thompson and Duncan, 1995) originating from magnetars where a major reconfiguration of the global magnetic field has occurred (Lyutikov, 2006). Magnetar flares also are loaded with mass entrained from the crust and the interior of the neutron star (Frail et al., 1999; Demidov and Lyubarsky, 2023; Patel et al., 2025). Thus, a complete view of giant flares requires modelling a mass-loaded magnetised explosion, reaching relativistic velocities. Here, ‘mass-loaded’ refers to finite plasma inertia and pressure within the Magnetohydrodynamical (MHD) framework, rather than explicit source terms in the continuity equation.
@@
The collection of an ample sample of fast radio bursts (FRBs), which are millisecond-duration transients of extragalactic origin, allows us to explore their diverse phenomenology, which indicates coherent emission from compact astrophysical sources (Lorimer et al., 2007; Petroff et al., 2019). Since their discovery, FRBs have challenged theoretical models, demanding mechanisms capable of releasing up to $10^{40}$ erg in radio waves over millisecond timescales, often with repetition, polarisation, and spectral properties indicative of strong magnetised environments (Marcote et al., 2020). The growing evidence linking at least a subset of FRBs to magnetars, particularly the association of FRB 200428 with the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), has reinforced the idea that magnetic energy release in highly magnetised neutron stars can power these events.
@@
Despite this association, the detailed physics linking magnetic activity to FRB emission remains uncertain. The proposed mechanisms span a broad range from magnetospheric reconnection and crustal cracking to relativistic shocks in magnetar winds (e.g. Lyubarsky, 2014; Beloborodov, 2017; Metzger et al., 2019). Most models assume magnetically dominated outflows capable of generating coherent radio emission through synchrotron maser, shocks, or antenna processes, and possibly a combination of these (Lyubarsky, 2021). Theoretical and observational developments suggest that magnetar magnetospheres may often be mass-loaded by plasma ejected during starquakes, bursts, or fallback episodes (Thompson and Duncan, 1995; Wadiasingh and Timokhin, 2019; Yuan et al., 2020). Moreover, the connection between magnetar bursts and flares and FRBs that has recently been confirmed (Bochenek et al., 2020; Tavani et al., 2021; Beniamini et al., 2025) stresses the necessity to consider mass-loaded magnetic detonations that might be related to an event originating from the neutron star interior and propagating to the magnetosphere (Bransgrove et al., 2025). The presence of dense plasma in a strongly magnetised environment can substantially modify the dynamics of magnetic explosions, their ability to accelerate particles, and the conditions for coherent emission.
@@
We explore mass-loaded magnetic explosions as potential origins of magnetar giant flares and FRBs. Using simplified relativistic magnetohydrodynamic (MHD) models, we examine how the interplay between the magnetic field, plasma mass loading, and self-similar expansion shapes the energetics and temporal properties of such events. The proposed structures can provide an appropriate environment that can lead to the generation of short, intense events with energetics and timescales consistent with observed FRBs.
@@
To model the coupled evolution of the magnetic field and plasma during these explosive events, we adopted the formalism developed in Gourgouliatos and Vlahakis (2010), which provides a self-consistent description of self-similar, co-expanding magnetised outflows. In this framework, the plasma and magnetic field evolve together under a homologous (Hubble-type) expansion, preserving the topology of the magnetic field while allowing for a time-dependent redistribution of electromagnetic and kinetic energy.
@@
The underlying concept traces back to the Prendergast (2005) family of self-similar, uniformly expanding, axisymmetric magnetic equilibria. These equilibria, later extended by Gourgouliatos and Lynden-Bell (2008) to include non-linear poloidal currents and toroidal fields, demonstrated that magnetic structures can expand coherently, a key insight for modelling transient magnetic explosions in compact objects. Building on this foundation, Gourgouliatos and Vlahakis (2010) formulated a relativistic configuration that couples the magnetic and velocity fields in a co-expanding MHD flow, enabling the analytical and numerical exploration of outflows that evolve self-similarly while conserving magnetic helicity and flux. In these models, the field originates from an appropriately twisted dipolar located at the origin. Due to the singularity at the origin, the field is studied beyond an inner radius that serves as a lower boundary condition.
@@
More recently, Barkov et al. (2022) applied similar ideas for numerical solutions of magnetic explosions in magnetically dominated and mass-loaded environments, demonstrating that magnetic energy release can drive relativistic outflows and shocks capable of producing bright, short-duration transients. In the same work, analytical solutions of force-free magnetic fields in the absence of thermal pressure and inertia forces were presented. These solutions do not diverge at the origin. It was found that the maximum expansion rate that can be attained by the expanding structure depends on the amount of twist of the magnetic field, with highly twisted magnetic fields not being able to expand at highly relativistic rates. In the absence of any thermal pressure and inertia, the pressure on the surface of the boundary of the spherical explosion is highly anisotropic due to the magnetic field structure, that is, in the dipole case, it is maximum at the equator and becomes zero at the axis of symmetry. When a structure like this is embedded within a uniform pressure environment, it will expand sideways on the equator and will form a highly anisotropic explosion, while it will lose its internal equilibrium and coherence. The inclusion of mass density and thermal pressure leads to configurations in the form of large-scale plasmoids whose surface pressure, including the electromagnetic term, is more isotropic, and under certain conditions, it can even become completely isotropic, with the surface magnetic field completely vanishing, as was shown in the non-relativistic approach (Gourgouliatos et al., 2010). These structures continue their uniform expansion when they are situated in a uniform density medium and maintain their internal equilibrium (Gourgouliatos and Lyutikov, 2012).
@@
In the present work, we extend these ideas by embedding mass-loaded magnetic explosions within the relativistically exapanding self-similar MHD formalism of Gourgouliatos and Vlahakis (2010). This allows us to consistently follow the co-evolution of magnetic fields and plasma during an explosive expansion and to explore the interplay between magnetic field strength, mass, and pressure and its effect on the expansion rate. This will allow us to identify the physical conditions under which these events can generate coherent radio emission with energetics and timescales comparable to those of magnetar flares and FRBs.
@@
This paper is organised as follows. In Sect. 2 we outline the physical and mathematical framework for magnetic plasmoid explosions and describe the key ingredients of our model. Section 3 presents the solutions and the exploration of the parameter space. The results are discussed in Sect. 4, focusing on energy release, plasma dynamics, and physical properties of the solution. In Sect. 5 we discuss the implications for astrophysical sources with a main focus on magnetar bursts and FRBs. We conclude in Sect. 6 and outline the limitations of the current model and the prospects for future work.
@@ 2 Mathematical formulation
We considered a fluid of rest-mass density $\rho$ and pressure $p$ containing a magnetic field $\bm{B}$. The electric and magnetic field obey Maxwell’s equations,
@@
where $j^{0}$ is the electric charge density, $\bm{j}$ is the electric current density, and $c$ is the speed of light. We further considered ideal MHD, and there is thus no electric field in the co-moving frame of the fluid. For an observer who sees the fluid moving at velocity $\bm{v}$, normalised to the speed of light, Ohm’s law becomes
@@
The continuity equation reads
@@
where $\Gamma$ is the Lorentz factor,
$\Gamma=\frac{1}{\sqrt{1-v^{2}}}.$ (7)
The momentum equation is
@@
where $p$ is the pressure, and the relativistic specific enthalpy is given by the following expression:
@@
where $\gamma$ is the ratio of the specific heats. Furthermore, the entropy equation reads
@@
In what follows, we assume a relativistic degenerate gas where $\gamma=4/3$. We further considered spherical coordinates $(r,\theta,\phi)$ and prescribed a velocity normalised to the speed of light given by the following expression:
@@
This velocity corresponds to uniform expansion from the centre. Thus, each fluid element moves at constant velocity without any acceleration, and the system remains implicitly at force balance. Here and throughout the paper, $v$ denotes the dimensionless radial expansion velocity $v=r/(ct)$.
It is possible to find self-similar, separable solutions using the polar angle $\theta$ and the expansion velocity $v$ as main variables. The full details of this procedure were shown by Prendergast (2005); Gourgouliatos and Lynden-Bell (2008); Gourgouliatos and Vlahakis (2010), and we summarise the basic steps for clarity and consistency below.
We express the magnetic field in terms of two functions: one related to the poloidal field $\Psi(\theta,v)$, and the other to the toroidal $T(\theta,v)$,
@@
Combining Ohm’s law, Eq. (5), with the momentum equation, Eq. (8), and taking the toroidal component, we deduce that the poloidal and toroidal functions are related to each other through the following expression:
@@
The pressure and the density are related through the entropy equation, Eq. (10), which is satisfied for
@@
where $Q(v,\theta)$ is an arbitrary function of $v$ and $\theta$. Upon integration of the continuity equation, Eq. (6), we obtain that
@@
and combining this with Eq. (14), we find that the pressure has the following form:
@@
where the quantities $Q$ $\tilde{p}$, $\tilde{\rho}$, all depending on the separation variables $v$ and $\theta$, are related to each other through the following expression:
@@
Next, we evaluate the current and charge densities through Maxwell’s equations, Eq. (2) and (4), and substitute them into the momentum equation, Eq. (8). We find that $\tilde{p}=\tilde{p}(\Psi)$ and that the system should satisfy the following equation:
@@
$+\frac{v^{2}}{(1-v^{2})^{2}}H\frac{dH}{d\Psi}+16\pi^{3}\sin^{2}\theta\,\frac{v^{4}}{(1-v^{2})^{3}}\frac{d\tilde{p}}{d\Psi}=0.$ (18)
@@
The function $g(v)$ determines the radial distribution of the poloidal magnetic flux under self-similar expansion. Its first non-zero root, $v_{0}$, defines the outer boundary of the magnetised plasmoid, beyond which all field components vanish or become disconnected from the inner region. The overall shape of $g(v)$ therefore controls both the spatial extent of the configuration and the presence or absence of surface currents and pressure anisotropies.
Upon integration, and considering the fact the $B_{\phi}$ component of the field must remain finite,
@@
with $c_{0}$ a non-negative constant, which governs the intensity of the toroidal field, or equivalently, the poloidal current. The pressure term becomes
@@
In the above expression, $\tilde{p}_{0}$ is the pressure on the axis and the spherical surface of the plasmoid; and it has to be a non-zero constant. There is no constraint on the sign of $c_{1}$, but the combination of $\tilde{p}_{0}$ and $c_{1}$ must ensure that $\tilde{p}$ remains non-negative throughout the solution.
Finally, substituting the expressions from Eq. (19) into Eq. (18), we obtain an ordinary differential equation for $g(v)$,
@@
The first three terms describe the poloidal magnetic structure under relativistic expansion, the fourth term represents the contribution of the toroidal magnetic field (poloidal current), and the final term accounts for pressure and inertia effects. The solution of the above equation provides the radial structure of the configuration.